OpenStudy (anonymous):

Solve 8x-5y=12 for y

OpenStudy (anonymous):

1) 8x-5y=12 -8x -8x Subtract 8x from both sides -5y= 12-8x Divide bothe sides by -5 so you isolate your y. -5y/-5= 12/-5 -8x/-5 y=12/-5+ 8x/5

OpenStudy (hoblos):

you need to solve for y first step subtract both sides by 8x you get -5y=12 - 8x then divide by -5, you get y = (12-8x)/(-5) = -12/5 + 8x/5